payton:

What is the solution to the system of equations? y = 1/3x - 10 2x + y = 4 (–8, 6) (–6, 16) (6, –8) (16, –6)

payton:

this one seems like the first one we did already...

ramen:

just try to do the same thing as the first problem we helped you with

ramen:

Yes!

payton:

so would the answer be D?

darkknight:

y = A system of equations. y equals StartFraction one-third EndFraction minus 10. 2 x plus y equals 4.x – 10? what is this supposed to be.

ramen:

i don't really understand what the first equation is bc it is weirldly typed

ramen:

but if you plug those numbers in and it works, than it is the right anwer.

ramen:

could you please retype the first equation plz?

ramen:

thanks

payton:

done

ramen:

just tell us how you got ur work then we can check it

payton:

ummmm so i combined -10 and -4

ramen:

actually ill see if what you got is right. hold up

payton:

??

ramen:

are you sure you typed the equations correctley for the problem?

payton:

yes

ramen:

you said ur answer was D the 4th one?

payton:

uhh??

ramen:

ok well ill take you step by step.

payton:

okay thanks

ramen:

1. since 2x+y=4 , the equation 1/3x-10 would be substituted into Y. this would be 2x+1/3x-10=4

ramen:

after this step, isolate the Xs to one side and the numbers without X to the other side.

ramen:

we would end up with 2x+1/3x=10+4

payton:

it showed me a graph for some reason

ramen:

oh. thats weird. but do you understand the steps so far?

payton:

2 1/3x -10 = 4

ramen:

that would be 2x+1/3x-10=4, then we move the 10 to the other side and would end up with 2x+1/3x=14. do you understand?

payton:

ok

ramen:

we combine like terms and get 2x+1/3x which would end up being 7/3x. 10+4=14. 7/3x=14. do you understand this?

payton:

kinda

ramen:

ok what do you need help understanding?

payton:

do i still have to divide 7/3x to equal 14

ramen:

Ill help you understand. So: 1. we start with y=1/3x-10 and 2x+y=4 2. Since 2x+y=4 has a Y inside the function, Y=1/3x-10 could be substituted in place of that Y

ramen:

7/3x=14 . To solve this you do 14 divided by (7/3) and that is your answer for X

ramen:

then to find Y, you would plug that X answer into all the X variables to find Y.

ramen:

hopefully you understand. Does this help?

payton:

so B?

ramen:

Does 14/(7/3) have a negative in it? No it doesn't so that means X can't be negative

payton:

c?

ramen:

Yes! correct good job

payton:

finally...lol

ramen:

do you need hlep understanding any of these concepts?

ramen:

sorry i was in a zoom

payton:

math is my weakest subject

ramen:

its alright.